# BioProBench: Comprehensive Dataset and Benchmark in Biological Protocol Understanding and Reasoning

### v1

@@ 4.1 Performance on Protocol Question Answering (PQA)
4.2 Performance on Step Ordering (ORD)
The ORD) task directly probes the LLMs’ ability to reconstruct the logical sequence of biological protocols. This task proved highly challenging for all evaluated models, highlighting significant limitations in inferring complex procedural dependencies (Table 3). Overall performance, measured by Exact Match (EM) accuracy requiring a perfectly ordered sequence, was low. The top-performing model, Gemini-2.5-pro-exp, achieved only 51.80% EM, indicating failure to capture the complete procedural flow in nearly half the cases. Other models lagged, with scores ranging down to 32.73% for GPT-4-turbo, reinforcing this bottleneck. However, Kendall’s Tau ($\tau$), measuring pairwise ranking correlation, presented a more positive view. $\tau$ scores were considerably higher, with Gemini-2.5-pro leading at 0.8104, and several others exceeding 0.73. This suggests models possess a reasonable grasp of local, pairwise step relationships even when failing to assemble the global sequence correctly.
@@ 4.4 Evaluating Reasoning Capabilities in Error Correction (REA-ERR)
Table 5: Performance Comparison on REA-ERR Task. The best value is highlighted in blue, and the runner-up value is highlighted in light blue.
@@
A salient observation arises from comparing REA-ERR with the standard ERR task. Strutured reasoning appeared to particularly enhance the performance of QwQ-32b and Deepseek-R1. This pronounced benefit is likely due to their training specifically focused on complex reasoning tasks. Notably, QwQ-32b’s F1 score surged from 59.41% in ERR to 71.40% in REA-ERR, rendering it the top F1 performer. Deepseek-R1 also maintained sencond-best F1 (69.17%) and best Consist.. This indicates that these models, trained with reasoning tasks, can effectively leverage explicit guidance to improve error identification. The REA-ERR results underscore that while structured reasoning process can enhance performance, particularly for models adept at reasoning, the consistency and validity of this elicited reasoning remain critical areas for advancement.
@@ 4.5 Performance on Protocol Generation (GEN)
As Table 6 shows, N-gram metrics are uniformly low: BLEU tops out at 10.23 (Qwen2.5-72b-instruct) and METEOR at 24.78 (Claude-3-7-sonnet), while ROUGE-L at just 20.70 (GPT-4-turbo). These results indicate limited lexical overlap and fluency, with METEOR’s slightly higher score reflecting modest gains in synonym and paraphrase handling. Domain-specific metrics reveal an even starker performance. The highest Keyword F1 is only 0.2457, suggesting models grasp some core terms but lack consistency. SR remains below 0.43, meaning over half of the essential steps are omitted or misrendered. SP scores between 0.20 and 0.32 reveal frequent inclusion of irrelevant or extraneous steps, highlighting low precision in generated steps. Among models, Claude-3-7-sonnet leads on SR and METEOR, Qwen2.5-72b-instruct excels in BLEU and keyword Recall, and GPT-4-turbo performs best on ROUGE-L and SP.
Notably, employing CoT prompting consistently degraded performance across most models and metrics in the GEN task. Claude-3-7-sonnet’s SR dropped from 0.4280 to 0.3918, and its METEOR score decreased slightly, while N-gram and Keyword scores saw more substantial reductions for most models under CoT. This suggests that untuned reasoning scaffolds can disrupt the generation of coherent, structured protocol text. To address this, BioProBench provides structured CoT exemplars, detailed in Section 2.2, that serve as a foundation for fine-tuning models to produce more fluent and logically consistent procedures.
@@ 6 Conclusion
Summary. In this work, we introduced BioProBench, the first large-scale, integrated multi-task benchmark for biological protocol understanding and reasoning. By curating 27K real-world protocols and generating over 556K structured task instances across five challenging tasks, alongside applying rigorous, domain-specific evaluation metrics, we exposed clear performance gaps in current large language models. Our evaluation exposed significant performance gaps in current LLMs, revealing struggles with deep temporal reasoning, causal logic, and generating safe, structured outputs, despite reasonable surface comprehension. By publicly releasing the dataset and evaluation suite, we offer a robust resource for advancing protocol-aware AI research.

### v2

@@ 4.2 Performance on Step Ordering (ORD)
The ORD task directly probes the LLMs’ ability to reconstruct the logical sequence of biological protocols. This task proved highly challenging for all evaluated models, highlighting significant limitations in inferring complex procedural dependencies (Figure 4b and Table 3). Overall performance, measured by Exact Match (EM) accuracy requiring a perfectly ordered sequence, was low. The top-performing model, Gemini-2.5-pro-exp, achieved only 51.80% EM, indicating failure to capture the complete procedural flow in nearly half the cases. Other models lagged, with scores ranging down to 32.73% for GPT-4-turbo, reinforcing this bottleneck. However, Kendall’s Tau ($\tau$), measuring pairwise ranking correlation, presented a more positive view. $\tau$ scores were considerably higher, with Gemini-2.5-pro leading at 0.8104, and several others exceeding 0.73. This suggests models possess a reasonable grasp of local, pairwise step relationships even when failing to assemble the global sequence correctly.
@@ 4.4 Evaluating Reasoning Capabilities in Error Correction (REA-ERR)
Table 5: Performance Comparison on REA-ERR Task. The best value is highlighted in blue, and the runner-up value is highlighted in light blue. Consist. unreported for domain-specific models as structured outputs absent.
@@
A salient observation arises from comparing REA-ERR with the standard ERR task: structured reasoning significantly boosts the performance of QwQ-32b and Deepseek-R1. QwQ-32b’s F1 score increased markedly from 59.41% to 71.40%, making it the top F1 performer. Deepseek-R1 also achieved the second-highest F1 (69.17%) and the best Consist. score. These results suggest that these models, trained on complex reasoning tasks, can effectively utilize explicit guidance to improve error identification. However, a notable finding from the REA-ERR task is the large gap between Consist. and Acc. scores. This discrepancy highlights that, although structured reasoning improves overall performance, ensuring the coherence and validity of the reasoning process itself remains a crucial challenge for future research.
@@ 4.5 Performance on Protocol Generation (GEN)
As Table 6 shows, N-gram metrics are uniformly low: BLEU tops out at 10.23 (Qwen2.5-72b-instruct) and METEOR at 24.78 (Claude-3-7-sonnet), while ROUGE-L at just 20.70 (GPT-4-turbo). These results indicate limited lexical overlap and fluency, with METEOR’s slightly higher score reflecting modest gains in synonym and paraphrase handling. Domain-specific metrics suggest a more pronounced performance gap. The highest Keyword F1 is only 0.2457, suggesting models grasp some core terms but lack consistency. SR remains below 0.43, meaning over half of the essential steps are omitted or misrendered. SP scores between 0.20 and 0.32 reveal frequent inclusion of irrelevant or extraneous steps, highlighting low precision in generated steps. Among models, Claude-3-7-sonnet leads on SR and METEOR, Qwen2.5-72b-instruct excels in BLEU and keyword Recall, and GPT-4-turbo performs best on ROUGE-L and SP.
Notably, employing CoT prompting consistently degraded performance across most models and metrics in the GEN task. Claude-3-7-sonnet’s SR dropped from 0.4280 to 0.3918, and its METEOR decreased slightly, while N-gram and Keyword scores saw more substantial reductions for most models under CoT. This suggests that untuned reasoning scaffolds can disrupt the generation of coherent, structured protocol text. To address this, BioProBench provides structured CoT exemplars, detailed in Section 2.2, that serve as a foundation for fine-tuning models to produce more fluent and logically consistent procedures. More experiments exploring the impact of CoT are in AppendixG.4.
@@ 6 Conclusion
Summary. In this work, we introduced BioProBench, the first large-scale, integrated multi-task benchmark for biological protocol understanding and reasoning. By curating 27K real-world protocols and generating over 556K structured task instances across five challenging tasks, alongside applying rigorous, domain-specific evaluation metrics, we exposed clear performance gaps in current large language models. Our evaluation exposed significant performance gaps in current LLMs, revealing struggles with deep temporal reasoning, causal logic, and generating safe, structured outputs, despite reasonable basic comprehension. By publicly releasing the dataset and evaluation suite, we offer a robust resource for advancing protocol-aware AI research.